# Revisiting Singlet Fermion Dark Matter with a Scalar Portal: Connecting Higgs Phenomenology and Strong Electroweak Phase Transition

### v2

@@ 5.3 Stochastic gravitational waves
As the temperature of the Universe drops below the critical temperature $T_{c}$, the electroweak-symmetric phase becomes metastable, and the transition to the true vacuum proceeds via thermal tunneling through the potential barrier. This process occurs through the nucleation of critical bubbles of the broken phase, whose probability is governed by the finite-temperature Euclidean (bounce) action Linde (1983); Affleck (1981); Grojean and Servant (2007). The bounce action exponentially suppresses the tunneling rate and determines the temperature at which the phase transition effectively completes.
@@ Appendix A Coefficients of $V_{\rm ct}(h,s)$
$\delta\lambda_{h}=-\frac{1}{2v^{3}}\left(v\frac{d^{2}V^{\rm CW}}{dh^{2}}-\frac{dV^{\rm CW}}{dh}\right)\Bigg|_{h=v,s=0}$ (73)
@@
$\delta\mu_{hs}=-\dfrac{1}{v}\frac{d^{2}V^{\rm CW}}{dhds}\Bigg|_{h=v,s=0}$ (75)

### v3

@@ 5.3 Stochastic gravitational waves
As the temperature of the Universe drops below the critical temperature $T_{c}$, the electroweak-symmetric phase becomes metastable, and the transition to the true vacuum proceeds via thermal tunneling through the potential barrier. This process occurs through the nucleation of critical bubbles of the broken phase, whose probability is governed by the finite-temperature Euclidean (bounce) action Linde (1983); Affleck (1981); Grojean and Servant (2007). The bounce action exponentially suppresses the tunneling rate and determines the temperature at which the phase transition effectively becomes complete.
@@ Appendix A Coefficients of $V_{\rm ct}(h,s)$
$\delta\lambda_{h}=-\frac{1}{2v^{3}}\left(v\frac{d^{2}V^{\rm CW}}{dh^{2}}-\frac{dV^{\rm CW}}{dh}\right)\Bigg|_{h=v,s=0},$ (73)
@@
$\delta\mu_{hs}=-\dfrac{1}{v}\frac{d^{2}V^{\rm CW}}{dhds}\Bigg|_{h=v,s=0},$ (75)